A quality control engineer at a potato chip company tests the bag filling machine by weighing bags of potato chips. Not every bag contains exactly the same weight. But if more than 16% of bags are over-filled then they stop production to fix the machine. They define over-filled to be more than 1 ounce above the weight on the package.

The engineer weighs 107 bags and finds that 22 of them are over-filled. He plans to test the hypotheses H0: p = 0.16 versus Ha: p > 0.16. What is the test statistic?

z =

(Please round your answer to two decimal places if necessary.)

( please be sure that answer is accurate, not 1 number off )
Answer #1

Solution:

Given

The null and alternative hypothesis is

H0 : p = 0.16

Ha : p > 0.16

р = x / n = 22/107 = 0.21

1 - P0 = 1 - 0.16 = 0.84

Test statistic = z

= р - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.21 - 0.16 / [\sqrt(0.16 * 0.84) / 107]

= 0.05 / 0.035441

Test Statistic (z) = 1.41